The diagonal of a rectangle is 25 cm and the length iss 10 cm more than twice the width. What is the perimeter of the rectangle?
Answer by mananth(16946)

let the width be x
the length is 2x+10
Diagonal = 25
..
they form a right triangle with diagonal as hypotenuse.
x^2+(2x+10)^2=25^2
x^2 + 4x^2+40x+100=625
5x^2+40x-525=0
x^2+8x-105=0
x^2+15x-7x-105=0
x(x+15)-7(x+15)=0
(x+15)(x-7)=0
x=7 cm the width
length = 2x+10
=2*7+10
=24 cm